OpenStudy (anonymous):

write a polynomail function of minmum degree with real coefficients whos zeros include those listed; 2,-12,5+2i

OpenStudy (dtan5457):

You need to put them all as (x-q) and multiply them all out. Just note that if 5+2i is a root, so is it's conjugate, 5-2i

OpenStudy (anonymous):

so i did X^2-10x+25=-4 10x+29=0 (x-2)(x+12)(x^2-10x+29) =f(X)=x^4-95x^2+530x-696

OpenStudy (dtan5457):

Yeah. I put the imaginary roots as (x-5+2i)(x-5-2i), multiplied them out, and got the same thing as you. as long as (x-2)(x+12)(x^2-10x+29) is correct, i think you have the right answer.

OpenStudy (anonymous):

ok thnx!